OpenStudy (anonymous):

]Identify the correct x intercept and y intercept of the equation x + 7y = -21. x intercept = (3, 0); y intercept = (0, 21) x intercept = (-3, 0); y intercept = (0, -21) x intercept = (21, 0); y intercept = (0, 3) x intercept = (-21, 0); y intercept = (0, -3)

OpenStudy (karatechopper):

for y int, make x=0 and solve for x int make y=0 and solve

OpenStudy (amistre64):

given the form: ax+by=c yint = c/b xint = c/a

OpenStudy (amistre64):

and since the zeros are in the right spots, you just gotta determine the number

OpenStudy (karatechopper):

i say the second one is the answer... click good asnwer for amistre and me and mark please!!

OpenStudy (amistre64):

really? since -21/1 = -21 id go number 4

OpenStudy (karatechopper):

have you solved for your x int yes amistre?

OpenStudy (amistre64):

x + 7y = -21 x = -21/1 y = -21/7

OpenStudy (karatechopper):

cuz the way u sloved...i understand what u did u made a big mistake...

OpenStudy (amistre64):

when x=0 ... y = -3 still

OpenStudy (karatechopper):

t o what i think

OpenStudy (karatechopper):

or let me think tht over..

OpenStudy (amistre64):

yint = (0,-3) still #4

OpenStudy (amistre64):

when y=0 x=-21, still xint = (-21,0) still

OpenStudy (karatechopper):

ah yes no my mistake i figured out what was wrong..yes i was wrong..i put a dif variable in..whoops my mistake king amistre u r totally correct

OpenStudy (amistre64):

.... im prone to error, so its good to dbl chk :)

OpenStudy (karatechopper):

more like tripple!

OpenStudy (amistre64):

this thing says I good answered you, hmm; musta been while I was typing lol

OpenStudy (amistre64):

oh well, cant give it twice ;)

OpenStudy (anonymous):

it the 2 one

OpenStudy (karatechopper):

hahaha,

OpenStudy (amistre64):

not unless you typoed the question

OpenStudy (karatechopper):

@louis i just proved that number 2 was wrong with amistre..its number 4

OpenStudy (amistre64):

oh well, theyll be back ...

OpenStudy (karatechopper):

hmm

OpenStudy (karatechopper):

hehy amistre do u have any experience with trignometry?

OpenStudy (anonymous):

yes choice 4 im sorry

OpenStudy (amistre64):

i now trig, occasionally

OpenStudy (anonymous):

i put it on reversed lol

OpenStudy (karatechopper):

@amistre can you teach me some basic trig?